# Two-scale Neural Networks for Partial Differential Equations with Small Parameters

### v1

@@ 3. Numerical results
We also compare the two-scale NN solutions to those from multi-level neural networks [1], known for their capability to address high-frequency issues. In the multi-level neural network method, a critical technique for addressing high-frequency issues involves scaling the equation’s source term with a factor. To be specific, at the initial neural network level, $u_{0}$ is solved from the equation with a scaling factor $\mu_{0}$: $R_{0}(x,u_{0})=\mu_{0}f-\mathcal{L}u_{0}=0$. The NN solution to this equation is denoted as $\tilde{u}_{0}$, resulting in the initial approximation $\tilde{u}=\tilde{u}_{0}/\mu_{0}$. In the subsequent $i$-th level ($i\geq 1$) of the neural network, $u_{i}$ is solved from $R_{i}(x,u_{i})=\mu_{i}R_{i-1}(x,\tilde{u}_{i-1})-\mathcal{L}u_{i}=0$. The NN solution for this level is denoted as $\tilde{u}_{i}$. The updated approximation is then obtained by summing the initial approximation with the cumulative corrections up to the $i$-th level, i.e., $\tilde{u}=\sum_{i=0}^{L}\frac{\tilde{u}_{i}}{\prod_{k=0}^{i}\mu_{k}}$. For more details of the algorithm of the multi-level neural networks, we refer the readers to Section 4 of [1].
We specifically investigate the case $(a_{1},a_{2})=(1,4)$ with the multi-level neural networks. The hyper-parameters used for multi-level NN are summarized in Table 2, which are suggested in [1] for 2D numerical examples. The networks are first trained with Adam, followed by refinement with L-BFGS. We note that [1] does not explicitly specify how to select the initial scaling factor $\mu_{0}$; nonetheless, we choose $\mu_{0}=\max\limits_{x\in\Omega}|f(x)|$. For the subsequent levels, according to [1], we choose $\mu_{i}\,(i\geq 1)$ to be the inverse of the amplitude of a rough approximation of $\tilde{u}_{i}$ by extreme learning machine [13] (where only the weights and biases of the output layer are updated) is of order 1. As a result, as shown in Figure 10 (left), we find that the approximation by the initial layer $\tilde{u}=\tilde{u}_{0}/\mu_{0}$ is comparable to the ones using our two-scale NN method (shown in Figure 9(d)). However, as shown in Figure 10 (right), the approximation with the correction by the first level $\tilde{u}=\tilde{u}_{0}/\mu_{0}+\tilde{u}_{1}/(\mu_{0}\mu_{1})$ does not improve the accuracy. Similar phenomena are observed for $(a_{1},a_{2})=(2,3)$. We also discover that the results of multi-level NN are highly sensitive to scaling factors. Nevertheless, this comparison underscores the critical role of scaling—whether applied to the equation or the neural networks—in addressing high-frequency issues.
@@
(f) full and truncated loss history
Figure 9. (b)(d)(e)(f) are numerical results with the two-scale NN for Example 3.5 a), when $k=4,a_{1}=1,a_{2}=4$ using $(x^{2}-1)(y^{2}-1)N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon},y,y/\sqrt{\epsilon},1/\sqrt{\epsilon})$, $\alpha=1,\alpha_{1}=0,N_{c}=22500,$ epoch=40000. We enforce $e_{rel}(x,y)=0$ on the nodes where the exact solution $u=0$. The relative $l^{2}$ error: $e_{l^{2}}^{rel}=5.2195\times 10^{-4}.$
@@
Figure 10. absolute error between exact solution and left: $\tilde{u}=\tilde{u}_{0}/\mu_{0}$; right: $\tilde{u}=\tilde{u}_{0}/\mu_{0}+\tilde{u}_{1}/(\mu_{0}\mu_{1})$ for Example 3.5 a) when $k=4,a_{1}=1,a_{2}=4$ using $(x^{2}-1)(y^{2}-1)N(x,y)$. $\alpha=1,\alpha_{1}=0$, $N_{c}=22500$, using multi-level neural networks with hyper-parameters specified in Table 2.
@@
(c) full and truncated loss history
(d) absolute error
(e) relative error
Figure 11. numerical results for Example 3.5 a), when $k=3,a_{1}=2,a_{2}=3$ using $(x^{2}-1)(y^{2}-1)N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon},y,y/\sqrt{\epsilon},1/\sqrt{\epsilon})$, $\alpha=1,\alpha_{1}=0,N_{c}=22500,$ epoch=40000. We enforce $e_{rel}(x,y)=0$ on the nodes where the exact solution $u=0$. The relative $l^{2}$ error: $e_{l^{2}}^{rel}=5.2195\times 10^{-4}.$
@@
(e) full and truncated loss history
@@
Figure 12. numerical results for Example 3.5 b) when $k=10$ using $(x^{2}-1)(y^{2}-1)N(x,x/{\sqrt{\epsilon}},y,1/{\sqrt{\epsilon}},y/{\sqrt{\epsilon}},1/{\sqrt{\epsilon}})$ with $k=1/{\sqrt{\epsilon}}$, $\alpha=1$, $\alpha_{1}=0,N_{c}=22497,$ epoch=80000. The relative $l^{2}$ error is: $e_{l^{2}}^{rel}=3.72\times 10^{-3}$.
@@ 4. Conclusion and discussion
In this work, we construct the two-scale neural networks by explicitly incorporating small parameters in partial differential equations into the architecture of feedforward neural networks. The construction enables solving problems with small parameters in a simple fashion, without modifying the formulations of PINNs or adding Fourier features in the networks. Extensive numerical tests illustrate that the two-scale neural network method is effective and provides reasonable accuracy in capturing features associated with large derivatives in solutions. These features include boundary layers, inner layers, and oscillations.
Notably, when a parameter in the differential equations becomes extremely small, the method struggles to accurately capture large derivatives. However, for problems with extremely small parameters, the proposed two-scale neural network method may be employed as a good initial guess for further training.
@@ Acknowledgement
We thank Mr. Ziad Aldirany from Polytechnique Montréal for providing the code for multi-level neural networks, which we utilized in comparing our results. GEK acknowledges support by the DOE SEA-CROGS project (DE-SC0023191) and the MURI-AFOSR project (FA9550-20-1-0358).

### v2

@@ 3. Numerical results
We also compare the two-scale NN solutions to those from multi-level neural networks ( MLNN )[1], known for their capability to address high-frequency issues. In the multi-level neural network method, a critical technique for addressing high-frequency issues involves scaling the equation’s source term with a factor. To be specific, at the initial neural network level, $u_{0}$ is solved from the equation with a scaling factor $\mu_{0}$: $R_{0}(x,u_{0})=\mu_{0}f-\mathcal{L}u_{0}=0$. The NN solution to this equation is denoted as $\tilde{u}_{0}$, resulting in the initial approximation $\tilde{u}=\tilde{u}_{0}/\mu_{0}$. In the subsequent $i$-th level ($i\geq 1$) of the neural network, $u_{i}$ is solved from $R_{i}(x,u_{i})=\mu_{i}R_{i-1}(x,\tilde{u}_{i-1})-\mathcal{L}u_{i}=0$. The NN solution for this level is denoted as $\tilde{u}_{i}$. The updated approximation is then obtained by summing the initial approximation with the cumulative corrections up to the $i$-th level, i.e., $\tilde{u}=\sum_{i=0}^{L}\frac{\tilde{u}_{i}}{\prod_{k=0}^{i}\mu_{k}}$. For more details of the algorithm of the multi-level neural networks, we refer the readers to Section 4 of [1].
We specifically investigate the case $(a_{1},a_{2})=(1,4)$ with the multi-level neural networks. The hyper-parameters used for multi-level NN are summarized in Table 4, which are suggested in [1] for 2D numerical examples. The networks are first trained with Adam, followed by refinement with L-BFGS.
It is recommended in [1] to set $\mu_{i},(i\geq 1)$ as the inverse of the amplitude of $u_{E,i}$. Here, $u_{E,i}$ is a rough approximation of the solution $\hat{u}_{i}$ to $\mathcal{L}\hat{u}_{i}-R_{i-1}=0$ by the extreme learning machine (ELM) [15]. For $\tilde{u}_{i}$, it is of the form:
(3.4) $\sin(\boldsymbol{\omega_{M}}\pi(1-x))\sin(\boldsymbol{\omega_{M}}\pi(1-y))N(\gamma(x),\gamma(y)),$
where the Fourier feature: $\gamma(x)=[\cos(\boldsymbol{\omega_{M}}x),\sin(\boldsymbol{\omega_{M}}x)],$ and ${\boldsymbol{\omega_{M}}}=(\omega_{1},...,\omega_{M}),$ $M$ is specified in Table 4.
We compare the two-scale NN solution to the MLNN’s initial and first levels of approximated solutions: $\tilde{u}_{0}/\mu_{0}$ and $\tilde{u}_{0}/\mu_{0}+\tilde{u}_{1}/\mu_{0}\mu_{1}$, with the hyper-parameters in Table 4. The two-scale NN results are obtained using the NN size of $(3,10,1)$, trained for 2500 epochs with Adam, followed by 200 epochs using L-BFGS. The MLNN results are obtained both without and with the Fourier feature in (3.4). We denote the absolute error between the exact solution and the $i$-th level approximated solution as $e_{i}=\left|u-\sum_{k=0}^{i}\left(\tilde{u}_{k}/\prod_{j=0}^{k}\mu_{j}\right)\right|$. The loss history is scaled by dividing $(\prod_{k=0}^{i}\mu_{k})^{2}$, where $i$ denotes the level at which the current loss function values are being calculated.
As we can see in Figure 12 (a) and (d) (e), the predictions by the two-scale NN and MLNN without Fourier feature provide reasonable accuracy in approximating the exact solution. The figure shows that the initial level prediction from the MLNN is on par with the two-scale NN when trained with the same number of epochs. The MLNN’s first-level correction refines the accuracy to $10^{-3}$. The results of the MLNN with the Fourier feature (3.4) shown in Figure 12 (h) indicate that adding the Fourier feature significantly improves the accuracy.
The results suggest that in addressing high-frequency issues, the two-scale NN does not necessarily surpass the approach of adding Fourier features or MLNN in accuracy; however, it offers a simpler solution by avoiding special treatments such as selecting wave numbers for Fourier features or scaling factors for MLNN. To further compare the two-scale NN method with MLNN, we consider two more 1D problems with one boundary layer, to be presented in Examples 3.6 and 3.7.
@@
(f) loss history
(g) errors history
Figure 11. (b)(d)(e)(f)(g) are numerical results with the two-scale NN for Example 3.5 a), when $k=4,a_{1}=1,a_{2}=4$ using $(x^{2}-1)(y^{2}-1)N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon},y,y/\sqrt{\epsilon},1/\sqrt{\epsilon})$, $\alpha=1,\alpha_{1}=0,N_{c}=22500,$ epochs=35000. We enforce $e_{rel}(x,y)=0$ on the nodes where the exact solution $u=0$. The relative $l^{2}$ error: $e_{l^{2}}^{rel}=4.1907\times 10^{-4}.$
@@
(a) absolute error of two-scale NN solution
(b) loss history
(c) errors history
(d) $e_{0}$ with MLNN
without Fourier feature
(e) $e_{1}$ with MLNN
without Fourier feature
(f) loss history
(g) errors history
(h) $e_{0}$ (left) and $e_{1}$ (right) with MLNN, adding Fourier feature
Figure 12. (a)-(c): Results for Example 3.5 a) with two-scale NN: $(x^{2}-1)(y^{2}-1)N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon},y,y/\sqrt{\epsilon},1\sqrt{\epsilon})$, when $k=4,a_{1}=1,a_{2}=4$, with $\alpha=1,\alpha_{1}=0$, $N_{c}=10000$.
(d)-(g): MLNN with NN: $(x^{2}-1)(y^{2}-1)N(x,y)$ ;
(h): MLNN with Fourier feature in (3.4). Hyper-parameters are specified in Table 4.
@@
(c) absolute error
(d) relative error
(e) loss history
(f) errors history
Figure 13. Numerical results for Example 3.5 a), when $k=3,a_{1}=2,a_{2}=3$ using $(x^{2}-1)(y^{2}-1)N(x,x/\sqrt{\epsilon},1/\sqrt{\epsilon},y,y/\sqrt{\epsilon},1/\sqrt{\epsilon})$, $\alpha=1,\alpha_{1}=0,N_{c}=22500,$ epochs=40000. We enforce $e_{rel}(x,y)=0$ on the nodes where the exact solution $u=0$. The relative $l^{2}$ error: $e_{l^{2}}^{rel}=5.5318\times 10^{-4}.$
@@
(e) loss history
@@
Figure 14. Numerical results for Example 3.5 b) when $k=10$ using $(x^{2}-1)(y^{2}-1)N(x,x/{\sqrt{\epsilon}},y,1/{\sqrt{\epsilon}},y/{\sqrt{\epsilon}},1/{\sqrt{\epsilon}})$ with $k=1/{\sqrt{\epsilon}}$, $\alpha=1$, $\alpha_{1}=0,N_{c}=22497,$ epochs=80000. The relative $l^{2}$ error is: $e_{l^{2}}^{rel}=3.41\times 10^{-3}$.
Example 3.6 (1D ODE with one boundary layer, comparison with multi-level neural networks).
$-\varepsilon u^{\prime\prime}+u^{\prime}=1,\quad\quad u(0)=0,\quad u(1)=0.$
The example is the same as Example 5.2 in [1]. We compare the results of solving this problem using the two-scale NN method and MLNN. To generate the MLNN results, we replicate the neural network configuration with the Fourier feature as described in Example 5.2 of [1]:
(3.6) $\sin\left(\boldsymbol{\omega_{M}}\pi(1-x)\right)N(\gamma(x)),$
where $\gamma(x)=[\cos(\boldsymbol{\omega_{M}}x),\sin(\boldsymbol{\omega_{M}}x)]$, ${\boldsymbol{\omega_{M}}}=(\omega_{1},...,\omega_{M})$. We denote the absolute error between the exact solution and the $i$-th level approximated solution as $e_{i}=\left|u-\sum_{k=0}^{i}\left(\tilde{u}_{k}/\prod_{j=0}^{k}\mu_{j}\right)\right|$.
The MLNN results are obtained with the hyper-parameters listed in Table 5 and presented in Figures 15 and 16. For the case $\epsilon=10^{-2}$, as shown in Figure 15, even though we use only one level of correction, the accuracy of the approximation achieves $10^{-7}$ in the absolute error. The loss and error histories demonstrate fast convergence and are consistent with Figure 17 in [1]. The accuracy can be further refined to $10^{-9}$ if three levels of corrections are applied. However, for the case $\epsilon=10^{-3}$, as shown in Figure 16, the MLNN results have not captured the exact solution, despite the employment of three levels of corrections.
As a comparison, we present the results using the two-scale NN method in Figure 17, with the NN size of $(3,20,20,20,20,1)$ and hyper-parameters in Table 1. When $\epsilon=10^{-2}$, as illustrated in Figure 17 (a) to (d), the absolute error reaches the magnitude of $10^{-4}$ after 35000 iterations. This is less precise compared to the MLNN results, where an accuracy of the order of $10^{-7}$ is achieved with just one level of correction. When $\epsilon=10^{-3}$, we solve the equation directly without employing the successive training strategy in Algorithm 1. As depicted in Figure 17 (e) to (h), the two-scale NN method successfully captures the key features of the exact solution. The overall accuracy is similar to that shown in Figure 2 for Example 3.1 (a similar 1D problem with one boundary layer).
In summary, this case study demonstrates that although the two-scale NN method may not necessarily surpass the MLNN method in terms of accuracy, it offers a simple approach to solving PDEs with small parameters. Furthermore, as the parameter in the PDE becomes smaller, the two-scale NN method can still provide reasonable accuracy with only a little or no special treatment.
Hyper-parameters $\tilde{u}_{0}$ $\tilde{u}_{1}$ $\tilde{u}_{2}$ $\tilde{u}_{3}$
# Hidden layers 1 1 1 1
Width of each hidden layer 5 10 20 20
# Adam iterations 4000 4000 4000 10000
# L-BFGS iterations 500 1000 2000 0
# Wave numbers $M$ 3 5 7 3
Table 5. Hyper-parameters used in Example 3.6 with multi-level neural networks [1]
(a) absolute error $e_{1}$
(b) loss history
(c) errors history
Figure 15. Numerical results for Example 3.6 when $\epsilon=10^{-2}$, using MLNN adding the Fourier feature in (3.6), with one level of correction.
(a) $\tilde{u}_{0}$
(b) $\tilde{u}_{1}$
(c) $\tilde{u}_{2}$
(d) $\tilde{u}_{3}$
(e) absolute error $e_{3}$
(f) loss history
(g) errors history
Figure 16. Numerical results for Example 3.6 when $\epsilon=10^{-3}$, using MLNN adding the Fourier feature in (3.6), with three levels of correction.
(a) exact and NN solutions for $\epsilon=10^{-2}$
(b) absolute error
(c) loss history
(d) errors history
(e) exact and NN solutions for $\epsilon=10^{-3}$
(f) absolute error
(g) loss history
(h) errors history
Figure 17. Numerical results for Example 3.6 using $N(x,(x-0.5)/\sqrt{\epsilon},1/\sqrt{\epsilon})$:
(a)-(d): when $\epsilon=10^{-2}$, with $\alpha=1$, $\alpha_{1}=0$, $N_{c}=200$, epochs=35000;
(e)-(h): when $\epsilon=10^{-3}$, with $\alpha=1$, $\alpha_{1}=0$, $N_{c}=450$, epochs=35000.
Example 3.7 (1D steady-state Burgers’ equation).
$u^{\prime\prime}-kuu^{\prime}=0,\quad 0<x<1,\quad u(0)=-1,\quad u(1)=-2/(k+2).$
This example, a nonlinear 1D ODE with a boundary layer, is the same as Example 7 in [1]. We solve this problem with the two-scale NN method with the NN size of $(3,20,20,20,20,1)$ and discuss the results alongside those presented in Example 7 of [1] using MLNN. Here, we take $\epsilon=1/k$.
When $k=8$ ($\epsilon=1/8$, a larger $\epsilon$), as shown in Figure 18 (a) to (d), the accuracy of the two-scale NN method reaches $10^{-5}$. In contrast, the MLNN results for this example, as reported in [1], achieve an accuracy of $10^{-13}$ with three levels of correction. Although the two-scale NN results do not reach the accuracy of those provided by MLNN, it is worth mentioning that the two-scale NN results are obtained without adding Fourier features and are optimized with only a first-order optimizer (Adam).
When $\epsilon$ is modestly small, such as $\epsilon=10^{-2}$ ($k=100$), as shown in Figure 18 (e) to (h), the two-scale NN method can directly handle it without any special treatment, achieving an accuracy of $10^{-5}$. As $\epsilon$ decreases further, for example to $\epsilon=10^{-3}$ ($k=1000$), the successive training strategy in Algorithm 1 can be utilized to enhance the training outcomes. As demonstrated in Figure 19, the two-scale NN method achieves an accuracy of $10^{-3}$ for $\epsilon=10^{-3}$ by employing Algorithm 1 with starting $\epsilon_{0}=10^{-2}$ and $\ell=2$.
Similar to our observations in Example 3.6, this case study shows that while the two-scale NN method may not necessarily exceed the accuracy of the MLNN method, it provides a simple approach to solving PDEs with small parameters. When the parameter in the PDE is modestly small, the two-scale NN method can deliver reasonable accuracy without any special treatment. For significantly smaller values of $\epsilon$, the successive training strategy can be employed to facilitate reasonably accurate outcomes.
(a) exact and NN solutions for $\epsilon=1/8$
(b) absolute error
(c) loss history
(d) errors history
(e) exact and NN solutions for $\epsilon=10^{-2}$
(f) absolute error
(g) loss history
(h) errors history
Figure 18. Numerical results for Example 3.7 using $N(x,(x-0.5)/\sqrt{\epsilon},1/\sqrt{\epsilon})$:
(a)-(d): when $\epsilon=1/8$ $(k=8)$ , with $\alpha=1$, $\alpha_{1}=0$, $N_{c}=200$, epochs=35000;
(e)-(h): when $\epsilon=10^{-2}$ $(k=100)$, with $\alpha=1000$, $\alpha_{1}=0$, $N_{c}=450$, epochs=35000.
(a) exact and NN solutions
(b) absolute error
Figure 19. Numerical results for Example 3.7 when $\epsilon=10^{-3}$ using $N(x,(x-0.5)/\sqrt{\epsilon},1/\sqrt{\epsilon})$ and Algorithm 1, with parameters specified in Table 6.
$\epsilon$ $10^{-2}$ (starting $\epsilon_{0}$) $5\times 10^{-3}$ $2.5\times 10^{-3}$ $1.25\times 10^{-3}$ $10^{-3}$
$\alpha$ 1000 1000 1000 10000 10000
$\alpha_{1}$ 0 0 0 00 00
$N_{c}$ 450 450 450 450 450
LR PC $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$
epochs 35000 35000 35000 35000 35000
Table 6. Parameters in the loss function (2.2) and hyper-parameters for the successive training for Example 3.7, LR is the abbreviation for learning rate, PC is the piecewise constant scheduler in Table 1.
@@ 4. Conclusion and discussion
In this work, we construct the two-scale neural networks by explicitly incorporating small parameters in PDEs into the architecture of feedforward neural networks. The construction enables solving problems with small parameters simply, without modifying the formulations of PINNs or adding Fourier features in the networks. To enhance the accuracy of neural network predictions, especially for PDEs with smaller parameters, a successive training strategy is introduced. Extensive numerical tests illustrate that the two-scale neural network method is effective and provides reasonable accuracy in capturing features associated with large derivatives in solutions. These features include boundary layers, inner layers, and oscillations.
Notably, when a parameter in the PDEs becomes extremely small, the method struggles to accurately capture large derivatives. However, for problems with extremely small parameters, the proposed two-scale neural network method may be employed as a good initial guess for further training.
@@ Acknowledgement
We thank Professor Serge Prudhomme and Mr. Ziad Aldirany from Polytechnique Montréal for providing the multi-level neural networks code for Examples 1 and 5.2 in their paper [1] and Example 3.5 a) in this work. We also thank Professor Houman Owhadi from the California Institute of Technology for helpful discussions. GEK acknowledges support from the DOE SEA-CROGS project (DE-SC0023191) and the MURI-AFOSR project (FA9550-20-1-0358).